SPECTROSCOPY AND DYNAMICS OF THE Ã ^{2}A′′ STATE OF VINYL RADICAL \underline{AARON MANN}, X. L. CHEN, VLADIMIR LOZOVSKY, C. BRADLEY MOORE, Department of Chemistry, The Ohio State University, Columbus, Ohio.

The vibrational structure of the Ã ^{2}A′′ state of vinyl radical has been investigated by photofragment detection. The subsequent dynamics from the Ã^{2}A′′ state were elucidated with ion imaging, revealing the partitioning of energy between the products, H and C_{2}H_{2}. Results will be discussed in comparison to recent theoretical predictions.